solve the system by addition. x+y=3 and x-y=1
Answer by checkley75(3666)

x+y=3
x-y=1
-----------------------
2x=4
x=4/2
x=2 answer.
2+y=3
y=3-2
y=1 answer.
proof
2-1=1
1=1